Problem 14-23 Calculating the Cost of Equity [LO1] Minder Industries stock has a beta of 1.20. The company just paid a dividend of $.50, and the dividends are expected to grow at 6 percent. The expected return on the market is 11 percent, and Treasury bills are yielding 6.2 percent. The most recent stock price for the company is $79. a. Calculate the cost of equity using the DCF method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. Calculate the cost of equity using the SML method. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g.. 32.16.) a. DCF method b. SML method
Problem 14-8 Calculating Cost of Debt [LO2] Jiminy's Cricket rate of 5 percent 4 years ago. The bond currently sells for 109 percent of its face The company's tax rate is 23 percent. The book value of the debt issue is $55 million. In addition, the company has a second debt issue on the market, a zero coupon bond with 7 years left to maturity; the book value of this issue is $45 million, and the bonds sell for 77 percent of par. Farm issued a bond with 25 years to maturity and a semiannual coupon value. a. what is the company's total book value of debt? (Enter your answer in dollars, not millions of dollars, e.g. 1,234,567.) millions of dollars, e.g. 1,234,567) calculations and enter your answer as a percent rounded to 2 decimal places, e.g.. b. What is the company's total market value of debt? (Enter your answer in dollars, not c. What is your best estimate of the aftertax cost of debt? (Do not round intermediate 3216.) a. Total book value b. Total market value c. Cost of debt
Answer #1

14-23).

a). DCF Method:

kE = [{D0 x (1 + g)} / P0] + g

= [{$0.50 x 1.06} / $79] + 0.06 = 0.0067 + 0.06 = 0.0667, or 6.67%

b). SML Method:

kE = rF + beta[E(rM) - rF]

= 6.2% + 1.2[11% - 6.2%] = 6.2% + 5.76% = 11.96%

14-8).

a). Total Book Value of Debt = Book Value of First Issue + Book Value of Second Issue

= $55 million + $45 million = $100 million, or $100,000,000

b). Total Market Value of Debt = Market Value of First Issue + Market Value of Second Issue

= [109% x $55 million] + [77% x $45 million]

= $59.95 million + $34.65 million = $94.60 million, or $94,600,000

c). To find the kD of the first issue, we need to put the following values in the financial calculator:

INPUT 21*2=42 -1,090 (5%/2)*1,000=25 1,000
TVM N I/Y PV PMT FV
OUTPUT 2.17

So, kD of the first issue = 2 x r = 2 x 2.17% = 4.34%

To find the kD of the second issue, we need to put the following values in the financial calculator:

INPUT 7*2=14 -770 0 1,000
TVM N I/Y PV PMT FV
OUTPUT 1.88

So, kD of the second issue = 2 x r = 2 x 1.88% = 3.77%

kD = [w(First issue) x kD(First issue)] + [w(Second issue) x kD(Second issue)]

= [(59.95/94.60) x 4.34%] + [(34.65/94.60) x 3.77%] = 2.75% + 1.38% = 4.13%

After-tax kD = kD x (1 - t) = 4.13% x (1 - 0.23) = 3.18%